# Low-Rank Adaptation for Foundation Models: A Comprehensive Review

### v1

@@ 3.4 Theoretical Foundations
While the practical advantages of LoRA are evident, understanding its underlying principles from a theoretical perspective is crucial. This section addresses key questions regarding its effectiveness, optimal rank selection, roles of update matrices, and induced behavioral changes in theoretical aspects.
@@
Although LoRA restricts updates to a low-rank subspace, it effectively targets the gradients most responsible for significant transformations in network behavior based on the Equation (7). By focusing on these critical gradients, LoRA preserves the model’s ability to generalize, ensuring that the network remains sensitive to essential input variations while being highly parameter-efficient.
@@
Zeng and Lee [72] conducted a comprehensive study on the expressive power of LoRA across different architectures. (i) For fully connected neural networks, that LoRA can adapt any model to accurately represent a smaller target model if the LoRA-rank satisfies:
@@
Q3: What are the roles of update matrices and B? Zhu et al. [46] provide a comprehensive analysis of the distinct roles played by matrices and in LoRA. Their work reveals an inherent asymmetry in these matrices, which has important implications for fine-tuning efficiency and model generalization.
The authors [46] demonstrate that primarily functions as a feature extractor from the input, while projects these features towards the desired output. This asymmetry suggests that fine-tuning alone can be more effective than fine-tuning . Notably, their analysis shows that a randomly initialized can perform nearly as well as a fine-tuned one, challenging the conventional practice of updating both matrices. Building on this insight, Zhu et al. derive generalization bounds for different LoRA variants using an information-theoretic framework. When fine-tuning only B, the generalization bound takes the form:
@@
LoRA updates attention matrices with low-rank matrices as , introducing controlled perturbations:
@@ 4 Frontiers
Building upon the technical foundations discussed above, which establish the core components and mechanisms of LoRA, this section explores frontier developments that extend the capabilities of LoRA in novel directions. These frontier developments leverage and combine its fundamental principles to enable new functionalities, tackle more complex tasks, and address challenges in model adaptation.
@@ 4.1.1 LoRA Composition
Optimization-based Composition. LoRAHub [75] leverages CMA-ES [76] gradient-free optimization to determine optimal coefficients for combining LoRA modules. Through few-shot learning, it autonomously selects and integrates modules for new tasks without requiring manual expertise or gradient computation. Similarly, LoRA-Flow[77] introduces dynamic fusion weights to adjust the impact of different LoRAs at each generation step, determined by a fusion gate with minimal parameters. This approach outperforms baselines with static task-level fusion weights across various
Retrieval-based Composition. LoraRetriever [78] implements dynamic retrieval and composition of LoRA modules based on input prompts. It first embeds task-specific LoRAs into a shared space using instruction fine-tuning on a subset of tasks, then retrieves relevant modules using cosine similarity. The framework supports both module fusion and mixture strategies while maintaining efficient batch processing.
@@ 4.1.2 Generalized Framework
Dual-branch Framework. Hydra [80] presents a more generalized formulation by integrating both parallel and sequential LoRA branches within the model. The parallel branch learns task-specific features, similar to standard LoRA, while the sequential branch linearly combines pre-trained features. This dual branch enables Hydra to capture both task-specific adaptations and leverage general pre-trained knowledge, offering a comprehensive adaptation mechanism that improves performance across tasks.
Multi-PEFT United Framework. GLoRA [81] further make a generalization by unifying various parameter-efficient fine-tuning methods beyond LoRA. It introduces trainable support tensors to scale and shift weights, features, and biases, effectively subsuming methods like LoRA, adapter tuning, and prompt tuning within a single framework. GLoRA employs evolutionary search to determine optimal layer-wise configurations of these tensors, which can take scalar, vector, or low-rank matrix forms. Through structural re-parameterization, GLoRA incurs no additional inference cost while providing greater flexibility than previous PEFT methods.
@@
| Hydra [80] | Dual Branch; Combines parallel branch (new features) and sequential branch (pre-trained features). |
@@
| XGBLoRA [66] | Gradient boosting framework with rank-1 LoRA adapters; combinations of multi-weak learners. |
@@ 4.1.4 Mixture of Experts with LoRA
Another important branch in the development of advanced LoRA architectures is the combination of LoRA with Mixture of Experts (MoE). MoE is a neural network architecture where multiple “expert” sub-networks specialize in different input patterns [82]. A gating mechanism routes inputs to the most appropriate experts, allowing the model to handle a wide range of tasks efficiently [83]. Given the input , the MoE model computes
@@
By integrating LoRA with MoE, models learn multiple pairs of low-rank matrices (LoRA experts) instead of a single pair, with a router determining the weights or selection of experts based on the inputs. During fine-tuning, the pre-trained LLM weights remain fixed, while the LoRA experts and the router are trained, leveraging the parameter efficiency of LoRA and the specialization capabilities of MoE. The typical framework is illustrated in Figure 6 (B).
@@
Zadouri et al. [84] introduced MoV and MoLoRA, aiming to achieve full fine-tuning performance while updating less than 1% of parameters and improving generalization to unseen tasks. MoV and MoLoRA utilize (IA)3 vectors and LoRA adapters as experts, respectively, employing a soft merging strategy where all experts contribute to the output weighted by router probabilities.
Building on these concepts, Luo et al.[85] proposed MoELoRA, treating LoRA modules as experts within an MoE framework. MoELoRA incorporates multiple LoRA experts and a gating network, employing top- routing and a load-balancing loss[86] to prevent convergence towards a limited set of experts. The application of contrastive learning among experts mitigates the random routing problem often observed in MoE models [87].
However, the fixed number of LoRA experts, e.g., MoELoRA [85], lacks flexibility and can be redundant due to representation collapse or learned routing policy overfitting [88]. To address this, Gao et al. [89] introduced MoLA, a layer-wise expert distribution approach that enables flexible assignment of LoRA experts across different Transformer layers. MoLA employs a top- routing mechanism to select relevant experts for each input. Beyond improving performance and parameter efficiency, MoLA exhibited promising continual learning capabilities due to its sparse expert activation, allowing the model to retain knowledge from previous domains while adapting to new ones.
@@
LoRAMoE [90] introduces multiple LoRA experts integrated via a router network, using a localized balancing constraint to encourage some experts to focus on leveraging world knowledge for downstream tasks. It employs a top- routing strategy, enabling the model to maintain world knowledge while improving performance on multiple tasks. MoRAL [91] uses question-answer pairs from unstructured text and combines the multi-tasking abilities of MoE with LoRA’s parameter efficiency. It employs a soft routing mechanism where all experts contribute to the output weighted by router probabilities. MoRAL maintains performance on previously seen tasks while adapting to new domains, addressing catastrophic forgetting.
@@
(3) Task-based Integration These methods tackle domain specificity and task interference challenges. Domain specificity arises when models trained on general-purpose data lack the specialized knowledge required for specific domains like medicine [92] or finance [95]. Task interference occurs when multiple tasks and their datasets compete during training, leading to degraded performance across tasks [94, 93, 95].
To address domain specificity, Liu et al. [92] proposed MOELoRA for multi-task medical applications. MOELoRA introduces multiple LoRA experts, each consisting of low-rank matrices, with a task-motivated gating function controlling each expert’s contribution based on task identities. This approach allows for task-specific learning while maintaining a shared knowledge base across tasks. Feng et al. [95] introduced MOA, an end-to-end parameter-efficient tuning method for multi-task learning. MOA first trains individual LoRA modules for different tasks, and then combines them using a sequence-level routing mechanism based on domain metadata, allowing the flexible combination of domain-specific LoRAs. Buehler et al. [96] presented X-LoRA, employing a dynamic MoE strategy with a deep layer-wise token-level approach. Starting with pre-trained LoRA adapters, X-LoRA uses a gating strategy leveraging hidden states to dynamically mix adapted layers. This allows the model to create novel combinations to solve tasks, demonstrating strong performance in scientific applications.
To solve task interference, Gou et al. [93] addressed task conflicts in vision-language instruction tuning with MoCLE. This method introduces an MoE architecture that activates task-customized parameters based on instruction clusters, employing a cluster-conditional routing strategy and incorporating a universal expert to improve generalization to novel instructions. Chen et al. [94] proposed LLaVA-MoLE to mitigate data conflicts in instruction fine-tuning of Multimodal LLMs. It introduces a sparse MoE design with multiple LoRA experts and employs a token-level routing strategy where each token is routed to the top-1 expert. This allows adaptive choices for tokens from different domains, effectively addressing data conflicts.
In addition, Tian et al. [97] proposed HydraLoRA, an asymmetric LoRA architecture that challenges the conventional symmetric expert structure in MoE-based approaches. Through empirical analysis, they discovered that in multi-task settings, matrix parameters from different LoRA heads tend to converge while matrix parameters remain distinct. Building on this observation, HydraLoRA introduces an architecture with a shared matrix across all tasks and multiple task-specific matrices, employing a trainable MoE router to automatically identify intrinsic components within the training data.
@@ 4.2 LoRA for Continual Learning
The parameter-efficient nature of LoRA allows for incrementally updating models on new tasks while mitigating catastrophic forgetting [98, 99]. Several key advantages motivate the use of LoRA for Continual Learning (CL): (1) reduced computational costs compared to full fine-tuning, (2) natural isolation of task-specific knowledge, and (3) flexible combination of task-specific adaptations. Existing LoRA-based continual learning methods can be broadly categorized into three approaches: regularization-based method, task arithmetic-based method, and ensemble-based techniques.
Regularization-based approaches that leverage parameter constraints on LoRA updates as the primary mechanism to prevent catastrophic forgetting, focusing on preserving critical model parameters. O-LoRA [98] addresses catastrophic forgetting by constraining new task updates to be orthogonal to the subspace of previous tasks. It leverages the insight that LoRA parameters effectively capture the gradient subspace of a task. O-LoRA incrementally learns new tasks in orthogonal subspaces while keeping previous LoRA parameters fixed. This approach allows for effective knowledge accumulation without interference. Online-LoRA [100] is a task-free online continual learning framework for Vision Transformers that addresses catastrophic forgetting without relying on rehearsal buffers. It combines weight regularization to protect important parameters and loss dynamics monitoring to detect distribution shifts, enabling real-time model adaptation while maintaining performance across changing data streams. It does not assume any task boundaries.
Task arithmetic-based approaches that leverage task vector arithmetic on LoRA parameters [101]. Chitale et.al [101] apply arithmetic operations on LoRA parameters to combine knowledge from multiple tasks. This approach trains separate LoRA modules for each task and then uses task vector addition to create a task-agnostic model. A key insight is that LoRA parameters create semantic ”task vectors” in weight space that can be manipulated algebraically.
Ensemble-based works maintain and combine multiple task-specific LoRA modules. CoLoR [99] maintain separate LoRA modules for each task and use an unsupervised approach to select the appropriate module at inference time. CoLoR trains task-specific LoRA modules sequentially and combines them using prototype-based task identification. This allows for the isolation of task knowledge while enabling flexible combinations. AM-LoRA [102] uses multiple task-specific LoRA modules combined with an attention mechanism to integrate knowledge from different tasks. The attention-based mixing strategy enables adaptive knowledge integration while preventing catastrophic forgetting between tasks.
@@ 4.3 LoRA for Unlearning
LoRA facilitates the targeted removal of specific knowledge from foundation models without necessitating extensive retraining. This section categorizes and examines methodologies employing LoRA for unlearning, focusing on three primary categories: modular decomposition methods, optimization-based approaches, and progressive unlearning pipeline.
@@
Optimization-Based Approaches. These approaches that primarily rely on optimization methods to target and remove specific knowledge. Cha et al. [106] propose a LoRA initialization technique that adapts the low-rank matrices based on the data, weighted by Fisher information to prioritize parameter adjustments crucial to removing targeted knowledge. In a different approach, Gundavarapu et al. [107] explore a gradient-based fine-tuning approach with LoRA to selectively unlearn harmful or unwanted information. Utilizing gradient ascent combined with low-rank LoRA updates, this method refines specific parts of the model’s knowledge.
@@ 4.4 LoRA for Federated Learning
Privacy and Security. Privacy and security and protection is paramount in federated learning. FedIT [109] and FFA-LoRA [110] establish fundamental frameworks that combine federated learning with instruction tuning for LLMs. These frameworks implement FedAvg privacy-preserving mechanisms, ensuring instruction data remains on local devices while only encrypted LoRA parameters are transmitted and aggregated at the central server. Taking a different approach, PrivateLoRA [111] transmits only activations between the central cloud and edge devices to ensure data locality. Advancing security further, Huang et al. [112] integrate model slicing with Trusted Execution Environment (TEE) to guard against malicious attacks, employing server-side TEE for latter model layers and Sparsification Parameter Fine-tuning (SPF) with LoRA, achieving both security and performance without requiring client-side TEE.
@@
Computation Efficiency. Despite significant advances in privacy and security, FL faces a fundamental trade-off between model expressiveness and computation efficiency. Research shows that to guarantee the ability to fit any target model, the rank of LoRA must meet a lower bound that scales with embedding size [113] as discussed in Section 3.4. However, implementing such high ranks in practice would incur prohibitive communication and computation costs, particularly challenging for resource-constrained devices [114, 115, 116]. To address this challenge, FedGBA [66] combines ensemble learning with rank-1 adaptations to enable highly efficient and expressive federated fine-tuning of FFM. Similarly, FFA-LoRA [110] reduces computational overhead by fixing randomly initialized non-zero matrices and fine-tuning only zero-initialized matrices.
@@
Heterogeneity Handling. Another significant challenge in FL is managing heterogeneous data, devices and models across clients. FL models often suffer performance degradation when data across clients is non-independently and identically distributed (non-IID). SLoRA [117] addresses this through data-driven initialization, while FedLoRA [118] implements a stacking-based aggregation method enabling the aggregation of LoRA adapters with different ranks across heterogeneous clients. Additionally, pFedLoRA [119] introduces a homogeneous small adapter to facilitate federated clients’ heterogeneous local model training through iterative training for global-local knowledge exchange. Wagner et al. [120] introduce trust-weighted gradient aggregation schemes for on-device fine-tuning of large language models with limited data, utilizing LoRA to reduce communication and outperforming FedAvg and local tuning methods, particularly in diverse data scenarios. The approach effectively addresses data heterogeneity and scarcity in local datasets. Cho et al. [121] propose heterogeneous LoRA, which enables efficient federated fine-tuning of small foundation models on resource-limited devices by combining high and low-rank adaptations across clients, achieving optimal performance with minimal communication while preventing overfitting.
@@
Personalization. Personalization in FL involves adapting the global model to individual clients through various techniques such as local fine-tuning, LoRA adaptation, or hybrid approaches. This process requires careful balancing of model performance, privacy protection, and system efficiency while considering each client’s unique data distribution and resource constraints. FedHLT [122] and FedLFC [123] exemplify this approach in Personalized Multilingual Modeling, combining low-rank adaptation with language family clustering to address FFM’s challenges of high communication costs and parameter interference, achieving improved performance with reduced overhead compared to baseline models. PER-PCS [124] is a framework that allows users to safely share and collaboratively assemble personalized LoRA pieces for large language models, achieving comparable performance to individual fine-tuning while significantly reducing computational costs and storage requirements.
@@
The integration of LoRA with FFM represents a significant advancement in distributed machine learning, offering innovative solutions for privacy-preserving model adaptation while balancing efficiency, heterogeneity, and personalization requirements. Through various frameworks that combine privacy mechanisms, efficient parameter transmission, and adaptive techniques, LoRA have made FL more practical for resource-constrained environments while maintaining model expressiveness and security. As edge computing and IoT applications continue to evolve, the synergy between LoRA and FFM promises to transform distributed machine learning by enabling efficient, secure, and personalized model deployment at scale in privacy-sensitive domains.

### v2

@@ 3.4 Theoretical Foundations
While the practical advantages of LoRA are evident, understanding its underlying principles from a theoretical perspective is crucial. This section addresses key questions regarding its effectiveness, optimal rank selection, the roles of update matrices, and the behavioral changes it induces.
@@
Although LoRA restricts updates to a low-rank subspace, it effectively targets the gradients most responsible for significant transformations in network behavior based on Equation (7). By focusing on these critical gradients, LoRA preserves the model’s ability to generalize, ensuring that the network remains sensitive to essential input variations while being highly parameter-efficient.
@@
Zeng and Lee [79] conducted a comprehensive study on the expressive power of LoRA across different architectures. (i) For fully connected neural networks, they show that LoRA can adapt any model to accurately represent a smaller target model if the LoRA rank satisfies:
@@
Q3: What are the roles of update matrices and ? Zhu et al. [47] provide a comprehensive analysis of the distinct roles played by matrices and in LoRA. Their work reveals an inherent asymmetry in these matrices, which has important implications for fine-tuning efficiency and model generalization.
The authors [47] demonstrate that primarily functions as a feature extractor from the input, while projects these features towards the desired output. This asymmetry suggests that fine-tuning alone can be more effective than fine-tuning . Notably, their analysis shows that a randomly initialized can perform nearly as well as a fine-tuned one, challenging the conventional practice of updating both matrices. Building on this insight, Zhu et al. derive generalization bounds for different LoRA variants using an information-theoretic framework. When fine-tuning only , the generalization bound takes the form:
@@
LoRA updates attention matrices with low-rank matrices , introducing controlled perturbations:
@@ 4 Frontiers
Building upon the technical foundations discussed above, which establish the core components and mechanisms of LoRA, this section explores frontier developments that extend the capabilities of LoRA in novel directions. These frontier developments leverage and combine their fundamental principles to enable new functionalities, tackle more complex tasks, and address challenges in model adaptation.
@@ 4.1.1 LoRA Composition
Optimization-based Composition. LoRAHub [82] leverages CMA-ES [83] gradient-free optimization to determine optimal coefficients for combining LoRA modules. Through few-shot learning, it autonomously selects and integrates modules for new tasks without requiring manual expertise or gradient computation. Similarly, LoRA-Flow [84] introduces dynamic fusion weights to adjust the impact of different LoRAs at each generation step, determined by a fusion gate with minimal parameters and outperforming baselines that use static task-level fusion weights.
Retrieval-based Composition. LoraRetriever [85] implements dynamic retrieval and composition of LoRA modules based on input prompts. It embeds task-specific LoRAs into a shared space using instruction fine-tuning on a subset of tasks, then retrieves relevant modules using cosine similarity. The framework supports both module fusion and mixture strategies while maintaining efficient batch processing. Similarly, by clustering and selecting task and domain data, Adapters Selector [86] locates the appropriate adapter in an adapter index according to the detected domain and task. Alternatively, ComLoRA [87] associates a learnable embedding vector with each LoRA adapter, trains a selector based on input prompts for dynamic selection, and merges the chosen adapter into the base model during inference to preserve low latency.
@@ 4.1.2 Generalized Framework
Dual-branch Framework. Hydra [89] presents a generalized formulation by integrating both parallel and sequential LoRA branches within the model. The parallel branch learns task-specific features, similar to standard LoRA, while the sequential branch linearly combines pre-trained features. This dual-branch design enables Hydra to capture both task-specific adaptations and leverage general pre-trained knowledge, offering a comprehensive adaptation mechanism that improves performance across tasks.
Multi-PEFT United Framework. GLoRA [90] further generalizes LoRA by unifying various parameter-efficient fine-tuning methods. It introduces trainable support tensors to scale and shift weights, features, and biases, effectively subsuming methods such as LoRA, adapter tuning, and prompt tuning within a single framework. GLoRA employs evolutionary search to determine optimal layer-wise configurations of these tensors, which can take scalar, vector, or low-rank matrix forms. Through structural re-parameterization, GLoRA incurs no additional inference cost while providing greater flexibility than previous PEFT methods.
@@
| Adapters Selector [86] | Retrieval-based composition; dynamically selects adapters during inference. |
| ComLoRA [87] | Retrieval-based composition; embeds LoRA adapters with learnable vectors and uses a selector to pick the winner. |
@@
| Hydra [89] | Dual-branch design; combines parallel (new features) and sequential (pre-trained features) adaptations. |
@@
| XGBLoRA [69] | Gradient boosting framework with rank-1 LoRA adapters; combines multiple weak learners. |
@@ 4.1.4 Mixture of Experts with LoRA
Another important advancement in LoRA architectures is combining LoRA with Mixture of Experts (MoE), where multiple “expert” sub-networks specialize in different input patterns [91]. A gating mechanism routes inputs to the most appropriate experts, allowing the model to handle a wide range of tasks efficiently [92]. Given the input , the MoE model computes
@@
By integrating LoRA with MoE, models learn multiple pairs of low-rank matrices (LoRA experts) instead of a single pair, with a router determining the weights or selection of experts based on inputs. During fine-tuning, the pre-trained LLM weights remain fixed, while the LoRA experts and the router are trained, leveraging the parameter efficiency of LoRA and the specialization capabilities of MoE. The typical framework is illustrated in Figure 6 (B).
@@
Zadouri et al. [93] introduce MoV and MoLoRA, which achieve full fine-tuning parity while updating fewer than 1% of parameters and improving generalization to unseen tasks. These methods treat (IA)3 vectors and LoRA adapters as experts, respectively, and employ soft merging—where all experts contribute to the output weighted by router probabilities.
Building on this, Luo et al. [94] propose MoELoRA, which integrates multiple LoRA experts within a Mixture-of-Experts (MoE) framework. It uses top- routing and a load-balancing loss [95] to avoid expert collapse, while contrastive learning among experts mitigates random routing issues [96].
However, the fixed number of LoRA experts—for example, in MoELoRA [94]—lacks flexibility and can be redundant due to representation collapse or overfitting in the routing policy [97]. To address this, Gao et al. [98] introduce MoLA, a layer-wise expert allocation strategy that dynamically assigns LoRA experts across Transformer layers via top- routing. Its sparse activation not only improves parameter efficiency but also supports continual learning by preserving knowledge from prior tasks.
More recently, MoKA [99] replaces LoRA’s low-rank decomposition with Kronecker products within a sparse MoE architecture, compressing parameters while boosting performance. RepLoRA [100] reparameterizes LoRA matrices using a lightweight MLP within an MoE framework, reducing the sample complexity for accurate low-rank estimation from exponential to polynomial. Finally, GOAT [101] closes the gap with full fine-tuning by adaptively integrating SVD-structured priors and aligning LoRA-MoE optimization with full fine-tuned MoE—enhancing performance without additional computational cost.
@@
LoRAMoE [102] integrates multiple LoRA experts via a router network with a localized balancing constraint, allowing experts to focus on world knowledge for downstream tasks. It uses a top- routing strategy to balance world knowledge and task performance. MoRAL [103] combines question-answer pairs from unstructured text with MoE’s multitasking and LoRA’s efficiency. It uses soft routing, where experts contribute based on router probabilities, adapting to new domains while preventing catastrophic forgetting.
@@
(3) Task-based Integration. These methods tackle domain specificity and task interference challenges. Domain specificity arises when models trained on general-purpose data lack the specialized knowledge required for specific domains like medicine [104] or finance [107]. Task interference occurs when multiple tasks and their datasets compete during training, leading to degraded performance across tasks [106, 105, 107].
To address domain specificity, Liu et al. [104] propose MOELoRA for multi-task medical applications, employing multiple LoRA experts with a task-aware gating function that routes inputs based on task identity—enabling task-specific adaptation while sharing a common knowledge base. Building on this paradigm, Feng et al. [107] introduce MOA, an end-to-end parameter-efficient method that first trains task-specific LoRAs and then combines them via sequence-level routing conditioned on domain metadata, facilitating flexible composition of domain-specialized adapters. Further advancing dynamic specialization, Buehler et al. [108] present X-LoRA, which adopts a layer-wise token-level MoE strategy: starting from pretrained LoRAs, it dynamically mixes adapted layers using gating networks driven by hidden states, enabling novel adapter combinations and strong performance in scientific domains.
To solve task interference, MoCLE [105] is an MoE architecture that activates task-specific parameters based on instruction clusters, using a cluster-conditional routing strategy and a universal expert to enhance generalization. LLaVA-MoLE [106] proposes a sparse MoE design with multiple LoRA experts and a token-level routing strategy, allowing adaptive routing for tokens from different domains to address data conflicts. HMoRA [109] improves multi-task learning with a hierarchical hybrid routing strategy and auxiliary loss function, enhancing task distinction and token-context capture. MoSLD [110] shares the upper projection matrix among experts and applies a dropout strategy to enhance generalization across tasks. SMoRA [111] uses dynamic rank-wise activation, treating each LoRA rank as an expert to improve knowledge sharing and mitigate task conflicts in multi-task scenarios.
In addition, Tian et al. [112] propose HydraLoRA, an asymmetric LoRA architecture that challenges the conventional symmetric expert structure in MoE-based approaches. Through empirical analysis, they discovered that in multi-task settings, matrix parameters from different LoRA heads tend to converge while matrix parameters remain distinct. Building on this observation, HydraLoRA introduces an architecture with a shared matrix across all tasks and multiple task-specific matrices, employing a trainable MoE router to automatically identify intrinsic components within the training data.
@@ 4.2 LoRA for Continual Learning
The parameter-efficient nature of LoRA allows for incrementally updating models on new tasks while mitigating catastrophic forgetting [113, 114]. Several key advantages motivate the use of LoRA for continual learning (CL): (1) reduced computational costs compared to full fine-tuning, (2) natural isolation of task-specific knowledge, and (3) flexible combination of task-specific adaptations. Existing LoRA-based continual learning methods can be broadly categorized into three approaches: regularization-based methods, task arithmetic-based methods, and ensemble-based techniques.
Regularization-based approaches leverage parameter constraints on LoRA updates as the primary mechanism to prevent catastrophic forgetting, focusing on preserving critical model parameters. O-LoRA [113] addresses this by constraining new task updates to be orthogonal to previous tasks’ subspaces, allowing incremental learning while preserving earlier tasks. GORP [115] combines full and low-rank parameters in a unified low-rank gradient subspace, mitigating forgetting and improving efficiency. Online-LoRA [116] enables task-free online continual learning for Vision Transformers, using weight regularization and real-time loss monitoring to maintain performance across changing data. TreeLoRA [117] optimizes task structures using hierarchical gradient similarity and sparse updates, enhancing performance in continual learning.
Task arithmetic-based approaches leverage task vector arithmetic on LoRA parameters. Chitale et al. [118] apply arithmetic operations on LoRA parameters to combine knowledge from multiple tasks. Their method trains separate LoRA modules for each task and then uses task vector addition to create a task-agnostic model. A key insight is that LoRA parameters create semantic “task vectors” in weight space that can be manipulated algebraically.
Ensemble-based works maintain and combine multiple task-specific LoRA modules. CoLoR [114] maintains separate LoRA modules for each task and uses an unsupervised approach to select the appropriate module at inference time. It trains task-specific LoRA modules sequentially and combines them using prototype-based task identification. This allows for the isolation of task knowledge while enabling flexible combinations. AM-LoRA [119] uses multiple task-specific LoRA modules combined with an attention mechanism to integrate knowledge from different tasks. The attention-based mixing strategy enables adaptive knowledge integration while preventing catastrophic forgetting between tasks.
@@ 4.3 LoRA for Unlearning
LoRA facilitates the targeted removal of specific knowledge from foundation models without necessitating extensive retraining. This section categorizes and examines methodologies employing LoRA for unlearning, focusing on three primary categories: modular decomposition methods, optimization-based approaches, and a progressive unlearning pipeline.
@@
Optimization-Based Approaches. These methods rely on optimization techniques to selectively remove or suppress specific knowledge from foundation models. Cha et al. [123] propose a data-driven LoRA initialization strategy that weights low-rank updates by Fisher information, prioritizing parameters most critical for erasing targeted knowledge. In contrast, Gundavarapu et al. [124] employ gradient ascent with low-rank LoRA updates to unlearn harmful or unwanted information, directly steering model outputs away from undesirable behaviors.
@@ 4.4 LoRA for Federated Learning
Privacy and Security. Privacy protection is paramount in federated learning. FedIT [126] and FFA-LoRA [127] establish frameworks that combine federated learning with instruction tuning for LLMs. These systems implement FedAvg privacy-preserving mechanisms, keeping instruction data on local devices while transmitting only encrypted LoRA parameters to the central server. Taking a different approach, PrivateLoRA [128] exchanges only activations between the central cloud and edge devices to preserve data locality. Advancing security further, Huang et al. [129] integrate model slicing with trusted execution environments (TEEs), employing server-side TEEs for later model layers and sparsification parameter fine-tuning (SPF) with LoRA to achieve both security and performance without requiring client-side TEEs.
@@
Computational Efficiency. Despite significant advances in privacy and security, FL faces a fundamental trade-off between model expressiveness and computational efficiency. Research shows that to guarantee the ability to fit any target model, the rank of LoRA must meet a lower bound that scales with embedding size [79] as discussed in Section 3.4. However, implementing high ranks incurs significant communication and computation costs, especially for resource-constrained devices [130, 131, 132]. To mitigate this, FedGBA [69] combines ensemble learning with rank-1 LoRA updates for expressive yet efficient federated fine-tuning. FFA-LoRA [127] reduces overhead by fixing randomly initialized non-zero matrices and fine-tuning only zero-initialized ones. More recently, FedEx-LoRA [133] achieves exact model updates with minimal cost by injecting a residual error term into the frozen pretrained weights, preserving both accuracy and efficiency.
@@
Heterogeneity Handling. A key challenge in FL is managing heterogeneous data, devices, and models across clients, particularly in non-independently and identically distributed (non-IID) scenarios. SLoRA [134] tackles this via data-driven LoRA initialization. FedLoRA [135] aggregates client adapters of varying ranks through a stacking-based method. pFedLoRA [136] introduces a lightweight adapter to facilitate iterative global-local knowledge exchange during heterogeneous local training. Wagner et al. [137] enhance fine-tuning robustness via trust-weighted gradient aggregation. HetLoRA [138] combines high- and low-rank adaptations to prevent overfitting and enable communication-efficient fine-tuning on resource-constrained devices. Recently, LoRA-A2 [139] alternates between freezing LoRA modules and adaptively selecting ranks to mitigate aggregation inconsistency caused by data heterogeneity, ensuring stable performance under high heterogeneity.
@@
Personalization. Personalization in FL adapts the global model to individual clients using techniques such as local fine-tuning, LoRA adaptation, or hybrid approaches, balancing performance, privacy, and efficiency based on each client’s data and resource constraints. FedHLT and FedLFC [140] combine low-rank adaptation with language-family clustering to address FFM challenges of high communication costs and parameter interference, achieving improved performance with reduced overhead compared to baselines. PER-PCS [141] allows users to safely share and collaboratively assemble personalized LoRA pieces for LLMs, reducing computation costs while maintaining high performance. FDLoRA [142] employs dual LoRA modules—one for capturing global knowledge and another for client-specific adaptation—enabling efficient personalization of large language models while reducing communication and computation costs. pFedMixF [143] addresses non-IID data, sequential tasks, and varying client capabilities through mixture-based frequency aggregation, balancing personalization and generalization while mitigating catastrophic forgetting in federated environments.
@@
The integration of LoRA with FFM advances distributed machine learning by providing privacy-preserving model adaptation while balancing efficiency, heterogeneity, and personalization. Combining privacy mechanisms, efficient parameter transmission, and adaptive techniques, LoRA makes FL practical for resource-constrained environments without compromising model security or expressiveness. As edge computing and IoT evolve, the synergy between LoRA and FFM promises to advance distributed machine learning by enabling efficient, secure, and personalized model deployment in privacy-sensitive domains.